OpenStudy (cloverracer):

-(5)^-1 -1/5 -5 5 1/5

OpenStudy (anonymous):

5^-1 = 1/5 so -5^-1 = -1/5

OpenStudy (cloverracer):

ohh okay that makes sense! thanks!

OpenStudy (anonymous):

yw